OpenStudy (anonymous):

Can someone help sorry to bother thanks.A reporter covering a naturalization ceremony wants to randomly select new citizens to interview. Out of the 200 new American citizens, 7 of them are from Ecuador, and 10 of them are from Kenya. What is the probability that the reporter will randomly select a person from Ecuador first and then a person from Kenya to interview?

OpenStudy (anonymous):

was my answer wrong?

OpenStudy (anonymous):

my computer just crashed and posted it again sorry

OpenStudy (anonymous):

then i will write again \[\frac{7}{200}\times \frac{10}{199}\]

OpenStudy (anonymous):

um I did that and it came out as a decimal is there another way?

OpenStudy (anonymous):

you mean you want the answer as a fraction? it is a probability so it has to be between 0 and 1, a decimal seems ok to me

OpenStudy (anonymous):

yea the problem was to be a fraction and I am really confusing myself with other ways to do that problem sorry to bother u

OpenStudy (anonymous):

7/3980

OpenStudy (anonymous):

may ask how did it thanks

OpenStudy (anonymous):

sorry :)

OpenStudy (anonymous):

sure no need to be sorry

OpenStudy (anonymous):

are you asking how i reduced the fraction, or are you asking where i got the original answer from?

OpenStudy (anonymous):

oh u reduced I get it now thanks :)

OpenStudy (anonymous):

yw